OpenStudy (anonymous):

Someone please help!! this problem is freaking hard :d 6^4√2-14^3√2+5^3√2+7^4√2 ????????

OpenStudy (anonymous):

\[6\sqrt[4]{2}-14\sqrt[3]{2}+5\sqrt[3]{2}+7\sqrt[4]{2}\]

OpenStudy (anonymous):

\( 1078 \sqrt{2}\ )

OpenStudy (anonymous):

is this the question? it is identical to \[6x-14y+5y+7x\] combine like terms

OpenStudy (anonymous):

I assumed the question as given.

OpenStudy (anonymous):

the "like terms" are \[6\sqrt[4]{2}+7\sqrt[4]{2}=13\sqrt[4]{2}\]and \[-14\sqrt[3]{2}+5\sqrt[3]{2}=-9\sqrt[3]{2}\]